# largesymbols"0D largesymbols"0D

### v1

@@ Abstract
We explore pretraining estimators for structural econometric models. The estimator is “pretrained” in the sense that the bulk of the computational cost and researcher effort occur during the construction of the estimator. Subsequent applications of the estimator to different datasets require little computational cost or researcher effort. The estimation leverages a neural net to recognize the structural model’s parameter from data patterns. As an initial trial, this paper builds a pretrained estimator for a sequential search model that is known to be difficult to estimate. We evaluate the pretrained estimator on 14 real datasets. The estimation takes seconds to run and shows high accuracy. We provide the estimator at pnnehome.github.io. More generally, pretrained, off-the-shelf estimators can make structural models more accessible to researchers and practitioners.
@@ 1 Introduction
Section 4 evaluates the pretrained NNE using 14 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes lots of effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the estimates by pretrained NNE with those by SMLE. Their estimates are similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
@@ 2.2 The data
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, one should first standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. It is simple to scale the estimate to make it interpretable with respect to the original, non-standardized attributes.
@@ 3.2.2 The parameter prior
where $\sigma$ is a constant, $\chi_{1}$ is chi distributed, and $\boldsymbol{\nu}$ is a unit vector with an uniform distribution on the unit sphere. An implication of this prior is $\lVert\boldsymbol{\beta}\rVert^{2}\sim\sigma^{2}\cdot\chi_{1}^{2}$, which is invariant to $d^{(prod)}$. We use the same type of prior for $\boldsymbol{\eta}$ and $\boldsymbol{\alpha}$.
@@ 4 Applications
We evaluate the pretrained NNE’s performance with 14 real datasets. In particular, we compare it to SMLE, which is the prevailing approach to estimate search models. Below, we first briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. We will consider both variants of SMLE (i.e., smoothed and GHK-based).
@@ 4.1 Datasets
Our 14 datasets come from five different settings (Expedia, Taobao, Trivago, JD, and an anonymous online retailer). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
@@
Table 2: Summary Statistics of 14 Datasets
@@
Taobao: iPad mini 10,000 3 5 0 1.2% 10.8%
Taobao: iPad 4 10,000 3 5 0 1.2% 9.6%
@@
Anonymous: laptop 10,000 6 0 0 5.9% 21.5%
Anonymous: vacuum 10,000 4 0 0 8.5% 19.0%
Anonymous: washer 10,000 5 0 0 7.8% 22.1%
@@
(1) Taobao: iPad mini
Our 1st dataset comes from Jiang et al. (2021) that study consumer searches and purchases of tablet computers on Taobao.com, China’s largest online retailer. The dataset focuses on the top-selling tablet iPad mini and its top 20 sellers. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include age, gender, past purchase volume, a weekend dummy, and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include a dummy of whether the seller and consumer come from the same geographical location, a dummy for higher price, and a dummy for higher seller rating. There are no advertising attributes. There are 104,189 search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
(2) Taobao: iPad 4
Our 2nd dataset also comes from Jiang et al. (2021) and includes the searches and purchases of iPad 4 on Taobao.com. IPad 4 was the second top selling tablet after iPad mini. We again use a subsample of 10,000 search sessions to ease the computational burden of SMLE.
@@
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. [^5] https://www.kaggle.com/competitions/expedia-personalized-sort/overview The data for the contest is publicly available on Kaggle.com. It has been used in several papers to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data includes multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
@@
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. [^7] https://connect.informs.org/msom/events/datadriven2020 The data was made available to INFORMS members in 2019. JD.com is China’s second largest business-to-consumer online retail platform. The data covers many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
@@
This dataset includes the app channel from the JD.com data. There are 112,640 search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
(10) Anonymous: laptop
A marketing company, REES46, provides data on consumer searches and purchases in a large anonymous online store. The data is available on Kaggle.com. [^8] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store The store sells products across many categories. This dataset focuses on laptops. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. To ease the computational burden of SMLE, we take a subsample of 10,000 search sessions.
(11) Anonymous: vacuum
(12) Anonymous: washer
@@
The RecSys Challenge includes search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.
@@
This dataset consists of the search sessions from the mobile channel in the RecSys Challenge by Trivago.
@@ 4.2 Monte Carlo studies
Table 3: Estimates in Monte Carlo Study for Taobao: iPad mini
@@
$R=50$ $\boldsymbol{\lambda}=(2,2)$
$\alpha_{0}$6.5 -6.502 (0.013) -7.190 (0.017) -5.860 (0.028) -7.291 (0.035) -2.613 (0.014)
$\eta_{0}$ 5.5 5.499 (0.004) 5.445 (0.005) 6.278 (0.009) 6.212 (0.010) 4.957 (0.007)
$\eta_{1}$0.2 -0.183 (0.011) -0.235 (0.018) -0.341 (0.019) -0.290 (0.017) -0.333 (0.019)
$\eta_{2}$0.1 -0.082 (0.008) -0.114 (0.014) -0.154 (0.014) -0.130 (0.013) -0.147 (0.014)
$\eta_{3}$ 0.1 0.097 (0.008) 0.117 (0.015) 0.154 (0.015) 0.129 (0.013) 0.150 (0.014)
$\eta_{4}$0.1 -0.132 (0.013) -0.112 (0.020) -0.223 (0.022) -0.183 (0.019) -0.222 (0.021)
$\eta_{5}$ 0.1 0.120 (0.010) 0.110 (0.019) 0.162 (0.018) 0.131 (0.016) 0.164 (0.018)
$\beta_{1}$ 0.5 0.513 (0.003) 0.490 (0.003) 0.448 (0.006) 0.382 (0.005) 0.432 (0.006)
$\beta_{2}$0.8 -0.800 (0.003) -0.786 (0.003) -0.611 (0.007) -0.531 (0.007) -0.600 (0.007)
$\beta_{3}$ 1.5 1.512 (0.003) 1.471 (0.003) 1.168 (0.009) 1.007 (0.008) 1.151 (0.009)
@@
Our first set of tests are Monte Carlo studies. Specifically, for each of the 14 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with those simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. The true $\boldsymbol{\theta}$ is set to be close to its estimate later in Section 4.3.
Figure 2: Monte Carlo Study With the Taobao: iPad mini Dataset
@@
We first examine in detail the results from the Taobao: iPad mini dataset, reported in Table 4.2. Column 1 shows the true value of $\boldsymbol{\theta}$. Column 2 shows the estimate by the pretrained NNE. Columns 3 - 6 show the estimates by SMLE, where $R$ denotes the number of simulations per consumer and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects two smoothing factors, one for search choices and one for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But here, doing so worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table includes three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. We actually have tried a grid of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(2,1)$ gives us the smallest root mean square error (RMSE) in this application.
@@
The upper plot of Figure 2 reports the RMSEs of the different estimators. The smoothed SMLE has the highest RMSEs. The pretrained NNE has the lowest RMSE, outperforming even the SMLE that uses GHK and the very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that this likelihood simulation remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
@@
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset has been subsampled from the original dataset to reduce computation burden (see Section 4.1). The SMLE still takes between 1 and 8 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). In relative terms, the pretrained NNE is 3-4 orders of magnitude faster. [^11] For Taobao: iPad mini, we have also run estimations on the original dataset with 104,189 search sessions. The SMLE (GHK-based, $R=1000$) uses about 80 hours while the pretrained NNE uses about 4 seconds. It is also worth noting that SMLE requires additional implementation costs, such as tuning smoothing factors, tuning the optimization routine, trying starting points. These tasks require running the estimation multiple times and they are not reflected in Figure 2. Furthermore, coding the SMLE is quite complex, especially with the GHK simulations.
@@
Next, we expand the examination to all 14 datasets. The results are reported in Table 4. Here, we omit the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the Taobao: iPad mini dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
@@
It is worth noting that the sizes of the 14 datasets here are not large (see Table 2). In fact, recall that six of the 14 datasets are subsampled from their respective original datasets to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much larger than those indicated by Table 4.
Table 4: Monte Carlo Studies Across 14 Datasets
@@
Taobao: iPad mini 0.279 (0.007) 0.815 (0.015) 0.521 (0.015)
Taobao: iPad 4 0.298 (0.010) 0.715 (0.015) 0.474 (0.013)
@@
Anonymous: laptop 0.094 (0.003) 0.795 (0.010) 0.453 (0.009)
Anonymous: vacuum 0.095 (0.003) 0.352 (0.005) 0.176 (0.005)
Anonymous: washer 0.179 (0.009) 0.614 (0.007) 0.361 (0.008)
@@
Taobao: iPad mini 0.02 72.77 457.63
Taobao: iPad 4 0.02 76.95 486.60
@@
Anonymous: laptop 0.02 132.72 630.06
Anonymous: vacuum 0.02 59.42 371.52
Anonymous: washer 0.02 126.48 552.85

### v2

@@ Abstract
We explore pretraining estimators for structural econometric models. The estimator is “pretrained” in the sense that the bulk of the computational cost and researcher effort occur during the construction of the estimator. Subsequent applications of the estimator to different datasets require little computational cost or researcher effort. The estimation leverages a neural net to recognize the structural model’s parameter from data patterns. As an initial trial, this paper builds a pretrained estimator for a sequential search model that is known to be difficult to estimate. We evaluate the pretrained estimator on 12 real datasets. The estimation takes seconds to run and shows high accuracy. We provide the estimator at pnnehome.github.io. More generally, pretrained, off-the-shelf estimators can make structural models more accessible to researchers and practitioners.
@@ 1 Introduction
Section 4 evaluates the pretrained NNE using 12 real datasets. We conduct two sets of tests. The first set consists of Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices under a “true” $\boldsymbol{\theta}$. Pretrained NNE is faster than SMLE by 3-4 orders of magnitude. Further, it recovers $\boldsymbol{\theta}$ with accuracy equal to or higher than SMLE. The result holds even when compared against the SMLE that uses a specialized simulation algorithm for search models that takes lots of effort to implement (Jiang et al. 2021). The second set of tests estimates the search model entirely from real data. Because the true $\boldsymbol{\theta}$ is unknown, we compare the estimates by pretrained NNE with those by SMLE. Their estimates are similar. The pretrained NNE is again faster by 3-4 orders of magnitude.
@@ 2.2 The data
Our pretrained NNE assumes that $\boldsymbol{x}_{ij}^{(prod)}$, $\boldsymbol{x}_{ij}^{(ads)}$ and $\boldsymbol{x}_{i}^{(cons)}$ are all standardized to have zero means and unit variances. The means and variances here are computed across $i$ and $j$, e.g., the mean for $\boldsymbol{x}_{ij}^{(prod)}$ is $\frac{1}{nJ}\sum_{i=1}^{n}\sum_{j=1}^{J}\boldsymbol{x}_{ij}^{(prod)}$. If the attributes are not standardized, we standardize them before using the pretrained NNE to estimate $\boldsymbol{\theta}$. We then rescale the estimate of $\boldsymbol{\theta}$ to make it interpretable with respect to the original, non-standardized attributes.
@@ 3.2.2 The parameter prior
where $\sigma$ is a constant, $\chi_{1}$ is chi distributed, and $\boldsymbol{\nu}$ is a unit vector with a uniform distribution on the unit sphere. An implication of this prior is $\lVert\boldsymbol{\beta}\rVert^{2}\sim\sigma^{2}\cdot\chi_{1}^{2}$, which is invariant to $d^{(prod)}$. We use the same type of prior for $\boldsymbol{\eta}$ and $\boldsymbol{\alpha}$.
@@ 4 Applications
We evaluate the pretrained NNE’s performance with 12 real datasets. In particular, we compare it to SMLE, which is the prevailing approach to estimate search models. Below, we first briefly describe each dataset. Then, we conduct Monte Carlo studies that estimate the search model on real-data attributes but simulated consumer choices. Finally, we estimate the search model on the real data entirely.
Before proceeding, we briefly describe existing approaches to estimate search models, for readers unfamiliar with the literature. Standard SMLE with accept-reject simulations is infeasible for search models because it requires prohibitive computational costs. A common workaround in literature is to smooth the simulated likelihood function (e.g., Honka and Chintagunta 2016, Ursu 2018). However, estimates are sensitive to smoothing factors and tuning these factors requires costly Monte Carlo experiments. More recently, Jiang et al. (2021) and Chung et al. (2023) extend the Geweke-Hajivassiliou-Keane (GHK) simulation algorithm to the likelihood of search models. The GHK-based SMLE does not require smoothing, but is much more complex to implement. We will consider both the smoothed SMLE and the GHK-based SMLE.
@@ 4.1 Datasets
Our 12 datasets come from four different settings (an anonymous online retailer, Expedia, JD, and Trivago). Table 2 provides summary statistics. We see that these datasets vary substantially in the number of search sessions, numbers of attributes, buy rate, and search rate. So they cover a fairly wide range of real-world scenarios.
@@
Table 2: Summary Statistics of 12 Datasets
@@
eCommerce: laptop 10,000 6 0 0 5.9% 21.5%
eCommerce: vacuum 10,000 4 0 0 8.5% 19.0%
eCommerce: washer 10,000 5 0 0 7.8% 22.1%
@@
(1) eCommerce: laptop
Our first dataset consists of consumer searches and purchases in an anonymous large online store, provided by the marketing company REES46. The data is available on Kaggle.com. [^5] https://www.kaggle.com/datasets/mkechinov/ecommerce-behavior-data-from-multi-category-store The store sells products across many categories. This dataset focuses on laptops. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price and five brand dummies: Asus, HP, Lenovo, MSI, and Xiaomi (the baseline brand is Acer). There are no consumer or advertising attributes. There are a very large number of search sessions. To ease the computational burden of SMLE, we take a subsample of 10,000 search sessions.
(2) eCommerce: vacuum
(3) eCommerce: washer
@@
The International Conference on Data Mining (ICDM 2013) organized a contest to predict consumer searches and bookings for hotels on Expedia. The data for the contest is publicly available on Kaggle.com. [^6] https://www.kaggle.com/competitions/expedia-personalized-sort/overview It has been used in several papers to study consumer search behaviors (Ursu 2018, Greminger 2022, Compiani et al. 2023). The data includes multiple travel destinations. “Destination 1” refers to the largest destination. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include star rating, review score, a chain indicator, location score, price, and promotion. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The advertising attribute $\boldsymbol{x}_{ij}^{(ads)}$ includes the ranking of the hotel.
@@
The 2020 MSOM Data Driven Research Challenge provided data on consumer experiences on JD.com. The data was made available to INFORMS members in 2019.888https://connect.informs.org/msom/events/datadriven2020 JD.com is China’s second largest business-to-consumer online retail platform. The data covers many aspects of consumer experiences on the platform, from website browsing to order fulfillment (see Shen et al. 2020 for details). We focus on consumer search and purchase activities. The consumer attributes $\boldsymbol{x}_{i}^{(cons)}$ include a weekend dummy and a working-hours dummy. The product attributes $\boldsymbol{x}_{ij}^{(prod)}$ include price, a dummy for first-party sellers, and two unnamed attributes. There are no advertising attributes.
@@
This dataset includes the app channel from the JD.com data. There are a large number of search sessions. To ease the computational burden of SMLE, we use a subsample of 10,000 search sessions.
@@
The Trivago data includes search sessions from both the desktop channel and mobile channel. This dataset focuses on the desktop channel.
@@
This dataset consists of the search sessions from the mobile channel in the Trivago data.
@@ 4.2 Monte Carlo studies
Table 3: Estimates in Monte Carlo Study with eCommerce: Laptop
@@
$R=50$ $\boldsymbol{\lambda}=(4,2)$
$\alpha_{0}$5.0 -5.009 (0.006) -5.779 (0.010) -4.696 (0.023) -3.347 (0.011) -2.767 (0.011)
$\eta_{0}$ 4.8 4.797 (0.002) 4.732 (0.004) 4.851 (0.009) 4.981 (0.005) 3.999 (0.006)
$\beta_{1}$0.6 -0.603 (0.002) -0.579 (0.003) -0.375 (0.005) -0.457 (0.005) -0.366 (0.005)
$\beta_{2}$0.5 -0.488 (0.002) -0.485 (0.003) -0.300 (0.005) -0.370 (0.005) -0.295 (0.005)
$\beta_{3}$0.3 -0.292 (0.002) -0.290 (0.003) -0.188 (0.005) -0.228 (0.006) -0.183 (0.005)
$\beta_{4}$1.0 -1.010 (0.003) -0.968 (0.003) -0.627 (0.005) -0.771 (0.005) -0.616 (0.005)
$\beta_{5}$0.4 -0.399 (0.003) -0.390 (0.003) -0.255 (0.005) -0.315 (0.006) -0.250 (0.005)
$\beta_{6}$0.5 -0.489 (0.003) -0.479 (0.003) -0.316 (0.006) -0.392 (0.006) -0.312 (0.005)
@@
Our first set of tests are Monte Carlo studies. Specifically, for each of the 12 real datasets, we: (i) keep the product, consumer, and advertising attributes, but (ii) replace the search and purchase choices with those simulated by the search model under a “true” $\boldsymbol{\theta}$. Then, we try to recover $\boldsymbol{\theta}$ from this dataset. Knowing the ground truth of $\boldsymbol{\theta}$ allows us to measure estimation accuracy. The true $\boldsymbol{\theta}$ is set to be close to its estimate later in Section 4.3.
Figure 2: Monte Carlo Study with eCommerce: Laptop
@@
We first closely examine the results from the eCommerce: laptop dataset, reported in Table 4.2. Column 1 shows the true value of $\boldsymbol{\theta}$. Column 2 shows the estimate by the pretrained NNE. Columns 3 - 6 show the estimates by SMLE. The $R$ denotes the number of simulations per consumer in SMLE and $R=50$ is typical in literature. For the smoothed SMLE, $\boldsymbol{\lambda}$ collects the smoothing factors for search choices and for purchase choices. [^10] Some applications use the same smoothing factor for search and purchase choices. But doing so here worsens RMSE substantially. Whether multiple smoothing factors are necessary is generally application-dependent. The table shows three settings of $\boldsymbol{\lambda}$ to illustrate the sensitivity of SMLE’s estimate to $\boldsymbol{\lambda}$. We actually have conducted an expensive grid search of $\boldsymbol{\lambda}$, and $\boldsymbol{\lambda}=(4,1)$ gives the smallest root mean square error (RMSE) in this application.
@@
The upper plot of Figure 2 reports the RMSEs of the different estimators. The pretrained NNE has the lowest RMSE, outperforming even the SMLE that uses GHK and a very large $R=1000$. Although we generally expect MLE to be asymptotically efficient, the performance of SMLE is affected by simulation errors. In search models, simulating the likelihood is particularly difficult because of the large number of possible search and purchase combinations. Our result here suggests that this likelihood simulation remains challenging despite workarounds such as smoothing and GHK. The result also shows that pretrained NNE can achieve high estimation accuracy despite being off-the-shelf.
@@
The lower plot of Figure 2 reports the computation time needed for a single estimation run. All the times are measured on the same hardware. Recall that this dataset is a subsample from a single product category of the original multi-category data (see Section 4.1). [^11] In the original data, there are 624 categories. The laptop category has 41,918 search sessions. Our dataset is a subsample of 10,000 search sessions. Even so, the SMLE requires between 1.5 and 10 hours. The pretrained NNE takes about 1 second (so the orange bar is barely visible). This represents a time saving of 3 to 4 orders of magnitude. In addition, note that implementing SMLE requires additional costs not reflected in Figure 2, such as tuning smoothing factors, tuning optimization routine, trying starting points. Each of these tasks requires multiple estimation runs. Moreover, Figure 2 does not reflect the coding cost of SMLE, which can be substantial with GHK.
@@
Next, we expand the examination to all 12 datasets. The results are reported in Table 4. The table omits the smoothed SMLE because, as we have seen, its estimates are sensitive to the smoothing factors $\boldsymbol{\lambda}$ (and the optimal $\boldsymbol{\lambda}$ is application-dependent). The main finding with Table 4 is qualitatively the same as what we have seen with the eCommerce: laptop dataset. The pretrained NNE achieves either comparable or better estimation accuracy at drastically lower computational costs.
@@
It is worth noting that the sizes of the 12 datasets here are not large (see Table 2). In fact, recall that each dataset focuses on only a subset (a single product category, travel destination, or sales channel) of its source. Further, four of the 12 datasets have been subsampled to reduce the computation burden. Because computation time scales with data size, the absolute time savings on large datasets will be much greater than those indicated by Table 4.
Table 4: Monte Carlo Studies Across 12 Datasets
@@
eCommerce: laptop 0.094 (0.003) 0.795 (0.010) 0.453 (0.009)
eCommerce: vacuum 0.095 (0.003) 0.352 (0.005) 0.176 (0.005)
eCommerce: washer 0.179 (0.009) 0.614 (0.007) 0.361 (0.008)
@@
eCommerce: laptop 0.02 132.72 630.06
eCommerce: vacuum 0.02 59.42 371.52
eCommerce: washer 0.02 126.48 552.85